The directions read: find a common solution for the system of equations -3y+12x=9 and x-y=0. I think the answer is (1,1) but i'm not sure if i'm correct. Please look over it and if it's correct just say i did it right, but if it's wrong then give me the answer and show how you got to that answer.thank you for your time.
Answer by oscargut(686)

it is right but i add a solution too
-3y+12x=9
x-y=0 so x=y so -3y+12y=9 so 9y=9 so y=1 then x=1